OpenStudy (steven513):

In the square, A represents the area of the square. Find the value of A such that A is equal to two times the side of the square.

OpenStudy (danjs):

Recall, the area of a square is the side length squared, call side length X, then Area A A = x^2

OpenStudy (danjs):

It also says the area is 2 times the side length... A = 2*x together this gives A = x^2 = 2x

OpenStudy (danjs):

so you can solve that for a value of x... x^2 = 2x x^2 - 2x = 0 x*(x - 2) = 0 x = 0 or x - 2 = 0